# Decoupling and Damping: Structurally-Regularized Gradient Matching for Multimodal Graph Condensation

### v1

@@ 3.5. The Proposed SR-GM Framework
To address this dual challenge, as shown in Figure 1, we propose Structurally-Regularized Gradient Matching (SR-GM), a framework that re-engineers the condensation process through two synergistic components: Gradient Decoupling and Structural Damping.
@@ 3.5.1. Parametrization of the Synthetic Graph
In SR-GM, the synthetic graph $\mathcal{S}$ is parameterized by its node features $\mathbf{X}_{\mathcal{S}}$ and the parameters $\Phi$ of a structure generator network $g_{\Phi}$. The node features $\mathbf{X}_{\mathcal{S}}$ are learned directly. Following GCond (jin2021graph), the adjacency matrix $\mathbf{A}_{\mathcal{S}}$ is not directly parameterized but is generated as a function of the node features: $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$.
@@ 3.5.2. Gradient Decoupling via Orthogonal Projection
To resolve the root cause of instability, we perform ”gradient surgery” on the feature gradients (liu2021conflict; yu2020gradient) of each synthetic node. This process is a prerequisite for computing the structural regularization term and is integrated into the outer loop’s meta-gradient calculation.
@@
If a conflict exists ($\langle\mathbf{g}_{i,\text{text}},\mathbf{g}_{i,\text{image}}\rangle<0$), we project each gradient onto the normal plane of the other to eliminate the conflicting components, following the PCGrad (yu2020gradient) methodology:
@@
This operation yields a decoupled feature gradient matrix $\mathbf{G^{\prime}}_{\mathcal{S}}$, where each row is the reconciled gradient $\mathbf{g^{\prime}}_{i}=\mathbf{g^{\prime}}_{i,\text{text}}+\mathbf{g^{\prime}}_{i,\text{image}}$.
@@ 3.5.3. Structural Damping via Gradient Field Regularization
To counteract the structural amplification of noise, we introduce a novel regularization term that operates directly on the decoupled gradient field. Unlike traditional graph regularization (yang2021rethinking) on node features, which GNNs already implicitly perform, our regularizer enforces smoothness on the optimization directions themselves. Let $\mathbf{G^{\prime}}_{\mathcal{S}}$ be the matrix of decoupled feature gradients from Equation (7). The structural regularization term is the Dirichlet energy of this gradient field:
@@ 3.5.4. Final Objective and Algorithm
where $\lambda$ is a hyperparameter balancing gradient matching accuracy with gradient field smoothness.The gradient matching loss $\mathcal{L}_{\text{gm}}=D\left(\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t}),\nabla_{\theta}\mathcal{L}_{\mathcal{T}}(\theta_{t})\right)$ follows the implementation in GCond (jin2021graph). SR-GM design fundamentally repurposes the graph structure from a source of instability into a tool for stabilization. The complete procedure, which correctly interleaves the outer-loop update of $\mathcal{S}$ and the inner-loop update of $\theta$, is detailed in Algorithm 1.
@@
9:   Generate synthetic adjacency matrix $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$ using Eq. (8)
@@
14:   Construct decoupled gradient matrix $\mathbf{G}^{\prime}_{\mathcal{S}}$ from $\mathbf{G}_{\mathcal{S}}$ using Eq. (10) (11)
@@
16:   Compute structure regularization loss $\mathcal{R}_{\text{struct}}$ using Eq. (12)
@@ 4. Experiment
To empirically evaluate the effectiveness and generalization capability of our proposed SR-GM framework, we conducted a comprehensive set of experiments. This section is structured as follows: We begin by detailing the experimental setup, including datasets, evaluation protocols, and competing methods. Subsequently, we compare SR-GM against a series of representative baselines to demonstrate its superior performance. To further validate the robustness and architecture-agnostic nature of our framework, we then investigate its effectiveness when integrated with various GNN architectures. Furthermore, we perform experiments on SR-GM’s training convergence and modality correlation, aiming to explore its training efficiency and the impact of inter-modal conflict on multimodal graph condensation. Finally, we conduct extensive experiments on key hyperparameters—specifically, the damping coefficient $\lambda$ and the condensation ratio rr—thereby providing a thorough validation of SR-GM’s effectiveness on multimodal graph data. All the experiments are conducted on an NVIDIA RTX 3090 GPU.
@@
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCond SR-GM $\triangle$(%) Original
@@ 4.1. Experimental Setup
We evaluate the condensation performance of the proposed framework on 4 multimodal datasets, i.e., Ele-fashion (ni2019justifying) and Goodreads-NC (wan2018item) for inductive setting and Amazon-Sports, Amazon-Cloth for transductive setting. Following (cai2024multimodal), we construct binary classification datasets for Amazon-Sports and Amazon-Cloth from the Amazon e-commerce platform [^1] https://www.amazon.com/; reviews with ratings $\geq$ 4 are labeled as positive, and those with ratings ¡ 2 as negative. Text and visual features for the Amazon-Sports and Amazon-Cloth datasets are encoded and aligned with ImageBind (girdhar2023imagebind), while CLIP (radford2021learning) is used for the Goodreads-NC and Ele-fashion datasets. The resulting features have dimensions of 1024 for Ele-fashion and Goodreads-NC, and 2048 for Amazon-Sports and Amazon-Cloth, and in all cases, the text and image modalities each constitute half of the total dimensionality. For the inductive setting, the test graph is not available during training. Dataset statistics are shown in Table 1.
@@
To demonstrate the effectiveness of SR-GM, we benchmark it against a comprehensive suite of baselines, which can be categorized into coreset selection and graph condensation methods. For coreset selection, we induce a subgraph from a representative node subset. We evaluate three selection strategies: Random, Herding (welling2009herding), and K-Center (sener2017active). For graph condensation, we employ three methods. GCond (jin2021graph) and MSGC (gao2023multiple) aim to preserve performance by matching GNN training gradients. In contrast, SGDD (yang2023does) focus on maintaining global graph properties by preserving the graph’s spectral characteristics and connectivity patterns.
@@
We evaluate the quality of the condensed graphs by their ability to preserve performance on the downstream task of inductive node classification. For each baseline and our proposed SR-GM, we first generate a condensed graph containing $rN$ nodes from the original training graph of $N$ nodes, where $r\in(0,1)$ is the condensation ratio. The evaluation follows a two-stage protocol. In the training stage, a standard GNN model is trained exclusively on the small, condensed graph. Crucially, the original full graph is entirely discarded and inaccessible during this stage. In the subsequent test stage, the GNN model trained on the condensed graph is applied to the original, unseen test graph to perform inference on the test nodes. The primary metric for evaluation is the resulting test accuracy. To ensure robust conclusions, we repeat all experiments 5 times and report the mean accuracy and its variance.
4.2. Comparison with Baselines
In this subsection, we test the performance of a 2-layer GCN on the condensed graphs, and compare the proposed SR-GM with baselines. Table 2 reports node classification performance. As shown in Table 2, our proposed SR-GM achieves performance that is comparable to, and often superior to, that of the baseline methods across all datasets and condensation ratios. We make the following observations:
@@ 4.2. Comparison with Baselines
Our proposed SR-GM significantly outperforms most baseline methods across all datasets and condensation ratios. For instance, on the Goodreads-NC dataset at a 0.025% ratio, SR-GM achieves 66.43% accuracy, creating a substantial performance margin of nearly 9 percentage points over the strongest baseline, GCond.
@@
Beyond superior accuracy, SR-GM demonstrates exceptional stability, especially in settings where other methods falter. As seen in Table 2, GCond exhibit unstable performance on datasets like Ele-fashion. SGDD and MSGC even fail on Goodreads-NC. In stark contrast, SR-GM consistently produces high-performing condensed graphs with low variance across all runs. We attribute this robustness to our optimization objective, which is designed to mitigate the gradient conflict between different modalities.
@@
As expected, the accuracy of most methods improves as the condensation ratio ($r$) increases, since a larger condensed graph can store more information. However, it is crucial to note that the performance gap between SR-GM and the baselines is most pronounced at lower condensation ratios. For example, on Ele-fashion, the accuracy improvement of SR-GM over the next-best baseline is larger at $r=0.1\%$ than at $r=0.5\%$.
Figure 2. Comparison of condensation efficiency between our proposed SR-GM and the baseline GCond on the Ele-fashion and Goodreads-NC datasets. We can observe that SR-GM achieves comparable or superior peak performance to GCond with a faster convergence speed.
4.3. Condensation Efficiency
In this section, we empirically evaluate the condensation efficiency of SR-GM against GCond. This is achieved by analyzing the relationship between the performance of a node classification model and the number of training epochs spent optimizing the condensed dataset. Figure 2 illustrates this relationship, with each curve representing the performance trajectory and terminating upon reaching its peak.
As shown in Figure 2, SR-GM converges faster by 220 and 50 epochs on the Ele-fashion and Goodreads-NC datasets, respectively. Results demonstrates that SR-GM performances on par with, or even superior to GCond while maintaining a comparable convergence speed.
Dataset Methods GCN MLP MMGCN GAT SAGE Avg.
Ele-fashion GCond 78.87 74.94 71.53 34.65 64.46 64.89
$r=0.1\%$ SR-GM 79.91 81.71 75.88 39.31 81.50 71.66
Goodreads-NC GCond 62.07 49.06 43.81 33.41 53.56 48.38
$r=0.05\%$ SR-GM 67.76 57.34 51.36 34.33 64.16 54.99
Table 4. Test accuracy (%) of SR-GM on the Ele-fashion datasets when using different backbone models for both condensation and evaluation. SAGE is an abbreviation for GraphSAGE. Original means training on original graph.
Backbone Ratio ($r$) SR-GM GCond Original
[…]
4.4. Cross-Architecture Generalizability
To evaluate the cross-architecture generalizability of our method, we condense a graph using SR-GM with a GCN backbone and then use this single condensed graph to train several different GNN architectures. As reported in Table 3, we test a diverse set of models, including GraphSAGE (hamilton2017inductive), GAT (velickovic2017graph), a multimodal GNN (MMGCN) (wei2019mmgcn), and a simple MLP (taud2017multilayer). The performance across these varied architectures demonstrates that the graphs condensed by SR-GM are not overfitted to the condensation model and possess good generalization capabilities.
4.5. Architectural Agnosticism of SR-GM
A key advantage of our SR-GM framework is its architectural agnosticism, meaning it is not tightly coupled to a specific GNN architecture. To demonstrate this flexibility, we conducted an experiment where we varied the underlying graph neural network used within the SR-GM condensation process.
We selected a diverse set of models for this evaluation, including popular GNNs such as GCN and GraphSAGE. Crucially, we also included a standard MLP as a stress test. The MLP is entirely graph-agnostic and processes only node features, allowing us to evaluate SR-GM’s ability to condense structural information even when the internal model does not perform message passing.
As shown in Table 4, we report the performance on the Ele-fashion dataset, where for each run, the same model (e.g., GCN) was used for both condensation and downstream evaluation. The results compellingly show that SR-GM consistently yields highly effective condensed graphs across all tested architectures, including the MLP. This strongly validates our claim that SR-GM’s effectiveness is not contingent on a particular GNN backbone, highlighting its robustness and wide applicability.
Figure 3. Test accuracy (%) for SR-GM and GCond using different input features: text-only, image-only, and fused text-image features on Ele-fashion and Goodreads-NC.
4.6. Ablation Study on the Impact of Multimodal Features
To investigate the impact of multimodal features on the performance of graph condensation, we conduct a comprehensive ablation study. In this experiment, we apply the baseline GCond method to graphs constructed with three different types of input features: text-only, image-only, and fused text-image features. We then compare its performance against our proposed SR-GM, which is implemented within the GCond framework to specifically address multimodal challenges. The node classification test accuracy on the Ele-fashion and Goodreads-NC datasets for these configurations is presented in Figure 3.
As illustrated in the figure, the introduction of fused multimodal data adversely affects the standard GCond method. We observe that the inherent conflict between different modalities leads to an average performance degradation of approximately 6% for GCond when compared to its performance on single-modality inputs. This highlights a significant challenge for existing gradient-matching-based condensation methods when applied to multimodal graphs. In contrast, our proposed SR-GM successfully mitigates this inter-modality gradient conflict during the condensation process. Consequently, SR-GM significantly outperforms GCond on the multimodal graph condensation task, achieving substantial accuracy improvements of 14% on Ele-fashion and 9% on Goodreads-NC. This result underscores the robustness and effectiveness of SR-GM in handling complex, multimodal graph data.
4.7. Parameter Analysis of Regularization Weight $\lambda$
A critical hyperparameter in SR-GM is the regularization weight, denoted by $\lambda$, which significantly influences the model’s performance. To investigate its impact on the graph condensation process, we conduct experiments on the Ele-fashion dataset under various settings of $\lambda$ and condensation ratio ($r$). The resulting test accuracies are visualized in a heatmap, as illustrated in Figure 4, where darker shades indicate higher accuracy.
From the results, we draw two primary observations. First, SR-GM demonstrates robust performance across most configurations, achieving high accuracy on the Ele-fashion dataset. A notable exception occurs at $r$ of 0.05%, where performance degrades for $\lambda$ values of 500 and 1000. Second, the findings suggest that while SR-GM is largely insensitive to high values of $\lambda$, an important correlation exists between $r$ and $\lambda$. Specifically, as $r$ increases (i.e., the number of nodes in the condensed graph grows), a correspondingly larger $\lambda$ is required to maintain optimal performance and prevent a significant decline in accuracy.
5. Conclusion
Appendix A APPENDIX
@@
Table 5. Inductive performance on Ele-fashion, Goodreads-NC and Transductive performance on Amazon-Sports, Amazon-Cloth using T5VIT encoding. Performance is reported as test accuracy (%) for inductive and AUC score for transductive. ”Original” denotes models trained on original graph. The best accuracies are bolded and the second best are underlined. $\triangle$(%) denotes the improvements of SR-GM upon the sub-optimal result.
@@
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCond SR-GM $\triangle$(%) Original
@@
Table 6. SR-GM Ablation Study Results. D: gradient decoupling, R: structure regularization. $\Delta_{1}$: (SR-GM) - (SR-GM w/o D), $\Delta_{2}$: (SR-GM w/o D) - (SR-GM w/o D+R).
Dataset Ratio ($r$) Method CLIP (Aligned) T5VIT (Unaligned) CLIP Increments T5VIT Increments

### v2

@@ 2.5. The Proposed SR-GM Framework
To address this dual challenge, as shown in Fig. 2, we propose Structurally-Regularized Gradient Matching (SR-GM), a framework that re-engineers the condensation process through two synergistic components: Gradient Decoupling and Structural Damping.
@@ 2.5.1. Parametrization of the Synthetic Graph
In SR-GM, the synthetic graph $\mathcal{S}$ is parameterized by its node features $\mathbf{X}_{\mathcal{S}}$ and the parameters $\Phi$ of a structure generator network $g_{\Phi}$. The node features $\mathbf{X}_{\mathcal{S}}$ are learned directly. Following GCond (Jin et al., 2021), the adjacency matrix $\mathbf{A}_{\mathcal{S}}$ is not directly parameterized but is generated as a function of the node features: $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$.
@@ 2.5.2. Gradient Decoupling via Orthogonal Projection
To resolve the root cause of instability, we perform ”gradient surgery” on the feature gradients (Liu et al., 2021; Yu et al., 2020) of each synthetic node. This process is a prerequisite for computing the structural regularization term and is integrated into the outer loop’s meta-gradient calculation.
@@
If a conflict exists ($\langle\mathbf{g}_{i,\text{text}},\mathbf{g}_{i,\text{image}}\rangle<0$), we project each gradient onto the normal plane of the other to eliminate the conflicting components, following the PCGrad (Yu et al., 2020) methodology:
@@
This operation yields a decoupled feature gradient matrix $\mathbf{G^{\prime}}_{\mathcal{S}}$, where each row is the reconciled gradient $\mathbf{g^{\prime}}_{i}=[\mathbf{g^{\prime}}_{i,\text{text}}\oplus\mathbf{g^{\prime}}_{i,\text{image}}]$.
@@ 2.5.3. Structural Damping via Gradient Field Regularization
To counteract the structural amplification of noise, we introduce a novel regularization term that operates directly on the decoupled gradient field. Unlike traditional graph regularization (Yang et al., 2021) on node features, which GNNs already implicitly perform, our regularizer enforces smoothness on the optimization directions themselves. The structural regularization term is the Dirichlet energy of this gradient field:
@@ 2.5.4. Final Objective and Algorithm
where $\lambda$ is a hyperparameter balancing gradient matching accuracy with gradient field smoothness.The gradient matching loss $\mathcal{L}_{\text{gm}}=D\left(\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t}),\nabla_{\theta}\mathcal{L}_{\mathcal{T}}(\theta_{t})\right)$ follows the implementation in GCond (Jin et al., 2021). SR-GM design fundamentally repurposes the graph structure from a source of instability into a tool for stabilization. The detailed algorithm of SR-GM is shown in Algorithm 1.
@@
9:   Generate synthetic adjacency matrix $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$ with Eq. (9)
@@
14:   Construct decoupled gradient matrix $\mathbf{G}^{\prime}_{\mathcal{S}}$ from $\mathbf{G}_{\mathcal{S}}$ with Eq. (11)(12)
@@
16:   Compute structural regularization loss $\mathcal{R}_{\text{struct}}$ with Eq. (13)
@@ 3. Experiment
To verify the effectiveness and robustness of the SR-GM framework, we design comprehensive experiments to explore the following research questions: Q1: Can SR-GM outperform the other graph condensation methods? Q2: How does SR-GM perform under scenarios where multimodal features are unaligned? Q3: Can SR-GM offer competitive condensation efficiency? Q4: How do gradient decoupling and structural regularization individually and synergistically impact the condensation quality? Q5: Can SR-RM leverage the complementary information from multimodal features? Q6: How do different downstream settings and hyperparameters affect SR-GM’s generalizability? Due to space constraints, additional experimental results are provided in the Appendix A. All the experiments are conducted on an NVIDIA RTX 3090 GPU.
@@
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCDM GCond SR-GM $\triangle$(%) Original
@@ 3.1. Experimental Setup
We evaluate the condensation performance of the proposed framework on 4 multimodal datasets, i.e., Ele-fashion (Ni et al., 2019) and Goodreads-NC (Wan and McAuley, 2018) for inductive setting and Amazon-Sports, Amazon-Cloth for transductive setting. Following (Cai et al., 2024), we construct binary classification datasets for Amazon-Sports and Amazon-Cloth from the Amazon e-commerce platform [^3] https://www.amazon.com/; reviews with ratings $\geq$ 4 are labeled as positive, and those with ratings ¡ 2 as negative. Text and visual features for the Amazon-Sports and Amazon-Cloth datasets are encoded and aligned with ImageBind (Girdhar et al., 2023), while CLIP (Radford et al., 2021) is used for the Goodreads-NC and Ele-fashion datasets. The resulting features have dimensions of 1024 for Ele-fashion and Goodreads-NC, and 2048 for Amazon-Sports and Amazon-Cloth, and in all cases, the text and image modalities each constitute half of the total dimensionality. For the inductive setting, the test graph is not available during training. Dataset statistics are shown in Table 1.
@@
We benchmark SR-GM against a comprehensive suite of baselines. These methods can be categorized into coreset selection and graph condensation methods. For coreset selection, we induce a subgraph from a representative node subset. We evaluate three selection strategies: Random, Herding (Welling, 2009), and K-Center (Sener and Savarese, 2017). For graph condensation, we employ four methods. GCDM (Liu et al., 2022) directly aligns the class distributions of the original and condensed graphs. GCond (Jin et al., 2021) and MSGC (Gao and Wu, 2023) aim to preserve performance by matching GNN training gradients. In contrast, SGDD (Yang et al., 2023) focus on maintaining global graph properties by preserving the graph’s spectral characteristics and connectivity patterns.
@@
We evaluate the quality of the condensed graphs by their ability to preserve performance on the downstream task of node classification. For each baseline and our proposed SR-GM, we first generate a condensed graph containing $rN$ nodes from the original training graph of $N$ nodes, where $r\in(0,1)$ is the condensation ratio. The evaluation follows a two-stage protocol. In the training stage, a standard GNN model is trained exclusively on the small, condensed graph. Crucially, the original full graph is entirely discarded and inaccessible during this stage. In the subsequent test stage, the GNN model trained on the condensed graph is applied to the original, unseen test graph to perform inference on the test nodes. The primary metric for evaluation is the resulting test accuracy. To ensure robust conclusions, we repeat all experiments 5 times and report the mean accuracy and its variance.
3.2. Comparison with Baselines (Q1)
In this subsection, we test the performance of a 2-layer GCN on the condensed graphs, and compare the proposed SR-GM with baselines. Table 2 reports node classification performance. As shown in Table 2, we make the following observations:
@@ 3.2. Comparison with Baselines (Q1)
SR-GM outperforms most baseline methods across all datasets and condensation ratios. Notably, on the Goodreads-NC dataset (0.025% ratio), SR-GM achieves 66.43% accuracy, surpassing the strongest baseline, GCond, by nearly 9 percentage points.
@@
Beyond accuracy, SR-GM demonstrates exceptional reliability. Unlike SGDD and MSGC, which fail to surpass basic coreset methods on Goodreads-NC (Table 2), SR-GM consistently maintains high performance with low variance.
@@
While increasing the condensation ratio typically boosts performance by expanding information capacity, it also makes optimization harder. Despite these challenges, SR-GM consistently outperforms other methods.
@@
Table 3. Inductive performance on Ele-fashion, Goodreads-NC and Transductive performance on Amazon-Sports, Amazon-Cloth using T5-ViT encoding. Performance is reported as test accuracy (%) for inductive and AUC score for transductive. ”Original” denotes models trained on original graph. The best accuracies are bolded and the second best are underlined. $\triangle$(%) denotes the improvements of SR-GM upon the sub-optimal result.
@@
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCDM GCond SR-GM $\triangle$(%) Original
@@
3.3. Analysis of Feature Alignment Effects on Multimodal Graph Condensation (Q2)
We conduct multimodal graph condensation experiments with unaligned feature encoding to simulate real-world scenarios with feature inconsistency, where textual features are encoded using T5 (Raffel et al., 2020) (512-dimensional) and visual features using ViT (Dosovitskiy, 2020) (512-dimensional). This heterogeneous encoding creates semantic gaps in the feature space, providing a challenging environment to validate SR-GM’s robustness. As shown in Table 3, SR-GM maintains superior performance across all datasets and condensation ratios under unaligned conditions, achieving improvements of 7.81% over the sub-optimal method on Goodreads-NC at 0.05% condensation ratio and 3.11% on Amazon-Sports at 0.3% condensation ratio, confirming the effectiveness of gradient decoupling in mitigating multimodal gradient conflicts.
Comparative analysis between Table 2 (aligned features) and Table 3 (unaligned features) reveals that aligned conditions yield higher original dataset accuracy, with Ele-fashion achieving 85.68% versus 84.47%, Goodreads-NC reaching 79.45% versus 76.93%, and Amazon-Cloth attaining 71.59% versus 68.31%, reinforcing the crucial role of feature alignment in multimodal graph learning. However, the radar charts in Fig. 4 demonstrate SR-GM’s distinctive advantage: in both aligned and unaligned scenarios, it exhibits the smallest performance gap relative to original datasets, with Fig. 4 particularly showing significantly smaller performance degradation compared to other condensation algorithms under unaligned conditions.
Table 4. Average training time per epoch (seconds) for GCond and SR-GM.
Dataset Ratio ($r$) GCond(s) SR-GM(s)
0.025% 8.36 9.40
Goodreads-NC 0.050% 8.55 9.42
0.100% 2.11 2.01
Ele-fashion 0.300% 2.30 2.33
0.500% 2.42 2.73
0.100% 2.06 2.18
Amazon-Sports 0.300% 2.13 2.21
0.500% 2.02 2.27
0.100% 2.10 2.50
Amazon-Cloth 0.300% 2.21 2.69
0.500% 2.35 2.92
Figure 5. Comparison of condensation efficiency between our proposed SR-GM and the baseline GCond on the Ele-fashion and Goodreads-NC datasets. We can observe that SR-GM achieves peak performance with a faster convergence speed.
Table 5. SR-GM Ablation Study Results. D: gradient decoupling, R: structural regularization. $\Delta_{1}$: (SR-GM) - (SR-GM w/o D), $\Delta_{2}$: (SR-GM w/o D) - (SR-GM w/o D+R).
Dataset Ratio ($r$) Method CLIP (Aligned) T5-ViT (Unaligned) CLIP Increments T5-ViT Increments